OpenStudy (anonymous):

Solve the following equation: x-6=7(x-3)

Parth (parthkohli):

Where are you having a problem?

OpenStudy (anonymous):

well im not sure what i do first

OpenStudy (anonymous):

PLZ @ParthKohli help him

OpenStudy (anonymous):

PLZ NOBODY ANSWER EXCEPT @ParthKohli

Parth (parthkohli):

You should start by distributing in the right-hand side.

Parth (parthkohli):

Have you heard of the distributive property?

OpenStudy (anonymous):

yes

Parth (parthkohli):

So can you distribute 7 into \(x - 3\)?

OpenStudy (anonymous):

im not sure

Parth (parthkohli):

Hint:\[7(x - 3) \implies 7(x) - 7(3) \]

OpenStudy (anonymous):

yes it can be

Parth (parthkohli):

So simplify \(7(x) - 7(3)\), please?

OpenStudy (anonymous):

be comes 7x-21

Parth (parthkohli):

Yep. Now, we have\[x - 6 = 7x - 21 \]Everybody hates Math, so we have to perform some magic to get \(x\) on one side and the rest of the friends on other.

Parth (parthkohli):

\(x\) is not liked by people, and we have three friends:\[x - 6 = 8\]We'll add 6 to both sides\[x - 6 + 6 = 8 +6 \]Because \(-6 + 6\) is zilch, we'd just cancel it out because it doesn't make sense to keep extra numbers when you could just remove them.\[ x = 8 + 6\]\[x = 14 \]

Parth (parthkohli):

That was just an example, anyway. What you have to do is get the \(x\) alone on one side.\[x - 6 = 7x - 21 \]Think of some possible ways you could do so.

Parth (parthkohli):

Hint: You may add 6 to both sides :)

OpenStudy (anonymous):

dont it become x=7x-15?

Parth (parthkohli):

Yeah, it does. Now what you gotta do is get the -15 alone too, because you want an answer like\[x = \text{some number} \]

Parth (parthkohli):

Can you subtract 7x from both sides?

OpenStudy (anonymous):

yea and you get 6x=-15 , i think

Parth (parthkohli):

You were almost right.\[x - 7x = -15 \]\[ -6x = -15\]Notice that negative sign there.

OpenStudy (anonymous):

ooh, i thought it was negative but i wasnt sure, but dont i divide -6 nd -15

Parth (parthkohli):

Yes, you do. But you can make it easier by cancelling the negative from both sides:\[\cancel{-}6x = \cancel-15\implies 6x = 15 \]

Parth (parthkohli):

Now you divide 6 from both sides to get \(x\) alone.

OpenStudy (anonymous):

which makes x=2.5

Parth (parthkohli):

Yes, that is right.

OpenStudy (anonymous):

thanks alot :D

Parth (parthkohli):

I have some things for you to note. Are you busy?

Parth (parthkohli):

Okay. Gotta have food. Will message you in some time :D